OpenStudy (anonymous):

let f(x)=8x^2+8x+2 Find f (x+h)-f(x)

OpenStudy (anonymous):

Plug x+h in : f ( x+h) = 8 ( x+h)² + ...?

OpenStudy (anonymous):

8(x+h)+2

OpenStudy (anonymous):

f ( x+h) = 8 ( x+h)² + 8(x+h)+2 Now expand it out:

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

( x+h)² = x² + 2hx + h² -> 8 ( x+h)² + 8(x+h)+2 = ....?